# A Kalman Filter for track reconstruction in very large time projection chambers

### v2

@@ Abstract
[inst1]organization=Univerity of Oxford, addressline=Wellington Square, city=Oxford, postcode=OX1 2JD, country=United Kingdom
@@ 2.1 The custom Kalman Filter
where $y$ is the vertical direction; $z$ is the drift direction; $\phi$ is the azymuthal angle of the transverse momentum i.e. the component of the momentum vector transverse to the drift direction; $\lambda$ is the “dip angle” between the transverse momentum and the total momentum vector; q is the charge sign of the particle and $p_{\text{T}}$ is the module of the transverse momentum. Note that the inverse transverse momentum can also be written in terms of track curvature $1/r$. The conversion is easily obtained using the standard formula for charged particles moving in a magnetic field:
@@
Figure 3: Flow chart describing the CKF algorithm. A seeding algorithm is used to obtain an estimate for the status of the system at the start of the trajectory $k=0$. Energy loss and multiple scattering corrections are applied to the estimate. The fit is then moved to the next point $k\rightarrow k+1$ either by applying the the BKF procedure or by using the mirror rotation, in the case that the limits of the $\sin\phi$ range have been surpassed. The algorithm is iterated point by point until the end of the trajectory is reached. Other minor modifications have been made to the CKF algorithm compared to the BKF in order to make the mirroring operation more stable.
The seeding strategy used for the CKF (as well as for the BKF) consists in a simple three-point circle finding algorithm and will be referred to from now on simply as Seed. In the plane perpendicular to the magnetic field, the trajectory of a charged particle is a circle. Since only one circumference will pass through any three points, one can find a point which is roughly at the start of the particle trajectory, one at the end and one in the middle and obtain the properties of the circle that passes though them. This equates to solving a system of three linear equations for three unknown variables: the coordinates of the center of the circumference $(x_{C},y_{C})$ and its radius $r$. From the circle properties and the coordinates of the three points, one can find an estimate for the state vector $s_{0}$ at the starting point $(x_{0},y_{0},z_{0})$: $y_{0}$ and $z_{0}$ are taken as the measured values; $q/p_{T_{0}}$ is converted form the track curvature $1/r$; $\sin\phi_{0}=x_{0}/r$; $\tan\lambda_{0}$ is estimated as the ratio between the displacement in the drift direction between the first and middle point and the correspondent circumference arch in the transverse plane. The three-point method for the estimation of the initial-state vector $s_{0}$ can be written as:
@@ 3 Toy Monte Carlo Simulation
To generate particle samples and validate the CKF algorithm, we employed a toy Monte Carlo (MC) tool called fastMCKalman [31]. This tool, stemming from the AliExternalTrackParam framework in the AliRoot code-base [28], was designed to be complemented by RootInteractive [32], an advanced statistical analysis tool. fastMCKalman has been developed with several objectives in mind: conducting rapid Monte Carlo (MC) simulations to evaluate tracking performance metrics, particle identification, and time-of-flight measurements across various detector setups. It was also designed to facilitate detailed studies on signal distortion in the ALICE detector and the derivation of performance metrics for its Run-3 upgrade and future iterations. In this discussion, we highlight the effectiveness of fastMCKalman in rapid simulation and tracking performance assessment within a TPC setting. It is important to note that this article represents the first use of this fast simulation tool and serves as part of its validation process. The use of fast Monte Carlo tools to complement more extensive generators such as Geant4 [40, 41, 42], are commonplace in high energy physics. The ALICE experiment, for example, implements independent Monte Carlo components in its simulation pipeline, to model detector effects that would be difficult to simulate using a more standard MC generator [43]. This is the envisioned future use of fastMCKalman. However, as of now, only basic effects such as multiple Coulomb scattering and energy loss through ionization have been implemented. In order to ensure the agreement of fastMCKalman with more traditional MC generators, the formulas used to simulate these effects are the same implemented by Geant4. Additionally the agreement of the fastMCKalman simulation with theoretical expectations is tested as part of the validation of the CKF algorithm in Sec. 3.2.
@@ 3.2 Tests and results: parameter scan sample
Figure 10: (a) CKF $q/p_{\text{T}}$ resolution $\sigma_{\text{CKF}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{CKF}}}$ as a function of the true $q/p_{\text{T}}$. (b) Ratio of the CKF $q/p_{\text{T}}$ resolution, over the theoretical expectations $\sigma_{\text{theo}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{theo}}}$, as a function of the true $q/p_{\text{T}}$. The histograms include all particles in the PS sample and are color-coded according to their particle types $t_{\textrm{ID}}=(0,1,2,3,4)=(\textrm{e},\mu,\pi,K,\textrm{p})$. Only tracks with a minimum of 10 points are considered. These plots have been produced using the interactive analytical tool ROOTInteractive [32]. The error bars are statistical. See Fig. 21 in Appendix for a zoom-in version.
@@
Figure 12: (a) CKF $q/p_{\text{T}}$ resolution $\sigma_{\text{CKF}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{CKF}}}$ as a function of the true $q/p_{\text{T}}$. (b) Ratio of the CKF $q/p_{\text{T}}$ resolution, over the theoretical expectations $\sigma_{\text{theo}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{theo}}}$, as a function of the true $q/p_{\text{T}}$. The histograms include all particles in the PS sample and are color-coded according to the radial resoulution $\sigma_{r\phi}$ used in the simulation. Only tracks with a minimum of 10 points are considered. These plots have been produced using the interactive analytical tool ROOTInteractive [32]. The error bars are statistical.
@@
The difference in performance due to the mirror rotation can be quantified using the ratios between the full reconstruction resolution $\sqrt{C_{44}^{\textrm{CKF}}}$ and the $\sqrt{C_{44}^{\text{BKF}}}$ obtained with the basic reconstruction at a given point along the track. Figure 15 shows the ratios $\sqrt{C_{44}^{\textrm{CKF}}}/\sqrt{C_{44}^{\text{BKF}}}$ as a function of the true $q/p_{\text{T}}$, color-coded according to the gas pressure $P_{\textrm{gas}}$. Figure 15(a) shows the results for a sample of only electrons, while Fig. 15(b) shows the results for a sample of muons and pions and Fig. 15(c) for a sample of protons and kaons. The points are again randomly taken along the reconstructed tracks and down-sampled to $10\%$ of the total. For the electron sample, an overall relative improvement of $\sim 60\%$ is shown at $p_{\text{T}}<100$ MeV/$c$, with peaks of up to $\sim 80\%$ for the lowest momentum tracks in low pressure environments. This behavior is in agreement with Eqs. 32 and 33 which show a dependency of the $1/p_{\text{T}}$ resolution on $1/\sqrt{N}$ and $1/\sqrt{L_{\textrm{Arm}}}$ respectively: using the mirroring technique, more space points of the tracks are used, resulting in larger $N$ and $L_{\textrm{Arm}}$. The dependence on the number of points is shown more clearly in Fig. 16, where the histograms are color coded for total number of points in the track, including those only accessible through the mirroring technique. Tracks containing more that 800 points are excluded for easier legibility of the results. It is shown that for the longest tracks, relative improvements of up to 80% can be achieved.
@@ 3.3 Tests and results: high-pressure sample
In Fig. 17 we show the relative momentum resolution and bias as a function of the true momentum for the three particle types present in the sample. At lower momenta ($p_{\text{true}}<1\text{GeV/}c$) the resolution is close to 2% for pions and muons while it is closer to 8% for the protons. In this momentum region the multiple scattering component of the resolution is dominant. This component is inversely proportional to the particles’ $\beta$ factor. With a given momentum, the proton, having a higher mass, will always have a smaller $\beta$ and thus a worse expected resolution. Furthermore, while muons and protons at lower energies have the chance to produce longer and even looping tracks inside the detector, protons will tend to loose their energy more quickly, again due to their masses. This is clearly shown in Fig. 18(b) where the average particle lengths as a function of their true momentum is shown. Somewhat significant biases are also shown at these lower momenta, especially for protons. For momenta $p_{\text{true}}>1\leavevmode\nobreak\ \text{GeV/}c$, the resolution is comparable for the three particle types and increases slowly with the particle momentum. At higher momenta the point resolution component is dominant, so a direct proportionality on the momentum is expected, with no distinction between the particle types. An inverse dependency on the lever arm and number of points and thus indirectly on the length is also expected, but as shown in Fig. 18(b), in this momentum range the average length of the track becomes roughly the same for all particle types.
@@ Acknowledgments
Figure 22: Relative momentum resolution (a) and bias (b) as function of of the tracks’ lever arm $L_{\textrm{Arm}}$ for the HP sample. The two properties are defined as $\mu$ and $\sigma$ of Gaussian fits of the momentum fractional residuals $p_{\text{reco}}/p_{\text{true}}-1$. The three particle types (protons, muons and pions) are drawn separately.

### v3

@@ Abstract
[inst1]organization=University of Oxford, addressline=Wellington Square, city=Oxford, postcode=OX1 2JD, country=United Kingdom
@@ 2.1 The custom Kalman Filter
where $y$ is the vertical direction; $z$ is the drift direction; $\phi$ is the azimuthal angle of the transverse momentum i.e. the component of the momentum vector transverse to the drift direction; $\lambda$ is the “dip angle” between the transverse momentum and the total momentum vector; q is the charge sign of the particle and $p_{\text{T}}$ is the module of the transverse momentum. Note that the inverse transverse momentum can also be written in terms of track curvature $1/r$. The conversion is easily obtained using the standard formula for charged particles moving in a magnetic field:
@@
Figure 3: Flow chart describing the CKF algorithm. A seeding algorithm is used to obtain an estimate for the status of the system at the start of the trajectory $k=0$. Energy loss and multiple scattering corrections are applied to the estimate. The fit is then moved to the next point $k\rightarrow k+1$ either by applying the BKF procedure or by using the mirror rotation, in the case that the limits of the $\sin\phi$ range have been surpassed. The algorithm is iterated point by point until the end of the trajectory is reached. Other minor modifications have been made to the CKF algorithm compared to the BKF in order to make the mirroring operation more stable.
The seeding strategy used for the CKF (as well as for the BKF) consists in a simple three-point circle finding algorithm and will be referred to from now on simply as Seed. In the plane perpendicular to the magnetic field, the trajectory of a charged particle is a circle. Since only one circumference will pass through any three points, one can find a point which is roughly at the start of the particle trajectory, one at the end and one in the middle and obtain the properties of the circle that passes though them. This equates to solving a system of three linear equations for three unknown variables: the coordinates of the center of the circumference $(x_{C},y_{C})$ and its radius $r$. From the circle properties and the coordinates of the three points, one can find an estimate for the state vector $s_{0}$ at the starting point $(x_{0},y_{0},z_{0})$: $y_{0}$ and $z_{0}$ are taken as the measured values; $q/p_{T_{0}}$ is converted from the track curvature $1/r$; $\sin\phi_{0}=x_{0}/r$; $\tan\lambda_{0}$ is estimated as the ratio between the displacement in the drift direction between the first and middle point and the correspondent circumference arch in the transverse plane. The three-point method for the estimation of the initial-state vector $s_{0}$ can be written as:
@@ 3 Toy Monte Carlo Simulation
To generate particle samples and validate the CKF algorithm, we employed a toy Monte Carlo (MC) tool called fastMCKalman [31]. This tool, stemming from the AliExternalTrackParam framework in the AliRoot code-base [28], was designed to be complemented by RootInteractive [32], an advanced statistical analysis tool. fastMCKalman has been developed with several objectives in mind: conducting rapid Monte Carlo (MC) simulations to evaluate tracking performance metrics, particle identification, and time-of-flight measurements across various detector setups. It was also designed to facilitate detailed studies on signal distortion in the ALICE detector and the derivation of performance metrics for its Run-3 upgrade and future iterations. In this discussion, we highlight the effectiveness of fastMCKalman in rapid simulation and tracking performance assessment within a TPC setting. It is important to note that this article represents the first use of this fast simulation tool and serves as part of its validation process. The use of fast Monte Carlo tools to complement more extensive generators such as Geant4 [40, 41, 42], is commonplace in high energy physics. The ALICE experiment, for example, implements independent Monte Carlo components in its simulation pipeline, to model detector effects that would be difficult to simulate using a more standard MC generator [43]. This is the envisioned future use of fastMCKalman. However, as of now, only basic effects such as multiple Coulomb scattering and energy loss through ionization have been implemented. In order to ensure the agreement of fastMCKalman with more traditional MC generators, the formulas used to simulate these effects are the same implemented by Geant4. Additionally the agreement of the fastMCKalman simulation with theoretical expectations is tested as part of the validation of the CKF algorithm in Sec. 3.2.
@@ 3.2 Tests and results: parameter scan sample
Figure 10: (a) CKF $q/p_{\text{T}}$ resolution $\sigma_{\text{CKF}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{CKF}}}$ as a function of the true $q/p_{\text{T}}$. (b) Ratio of the CKF $q/p_{\text{T}}$ resolution, over the theoretical expectations $\sigma_{\text{theo}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{theo}}}$, as a function of the true $q/p_{\text{T}}$. The histograms include all particles in the PS sample and are color-coded according to their particle types $t_{\textrm{ID}}=(0,1,2,3,4)=(\textrm{e},\mu,\pi,K,\textrm{p})$. Only tracks with a minimum of 10 points are considered. These plots have been produced using the interactive analytical tool ROOTInteractive [32]. The error bars are statistical. See Fig. 21 in Appendix for a zoom-out version.
@@
Figure 12: (a) CKF $q/p_{\text{T}}$ resolution $\sigma_{\text{CKF}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{CKF}}}$ as a function of the true $q/p_{\text{T}}$. (b) Ratio of the CKF $q/p_{\text{T}}$ resolution, over the theoretical expectations $\sigma_{\text{theo}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{theo}}}$, as a function of the true $q/p_{\text{T}}$. The histograms include all particles in the PS sample and are color-coded according to the radial resolution $\sigma_{r\phi}$ used in the simulation. Only tracks with a minimum of 10 points are considered. These plots have been produced using the interactive analytical tool ROOTInteractive [32]. The error bars are statistical.
@@
The difference in performance due to the mirror rotation can be quantified using the ratios between the full reconstruction resolution $\sqrt{C_{44}^{\textrm{CKF}}}$ and the $\sqrt{C_{44}^{\text{BKF}}}$ obtained with the basic reconstruction at a given point along the track. Figure 15 shows the ratios $\sqrt{C_{44}^{\textrm{CKF}}}/\sqrt{C_{44}^{\text{BKF}}}$ as a function of the true $q/p_{\text{T}}$, color-coded according to the gas pressure $P_{\textrm{gas}}$. Figure 15(a) shows the results for a sample of only electrons, while Fig. 15(b) shows the results for a sample of muons and pions and Fig. 15(c) for a sample of protons and kaons. The points are again randomly taken along the reconstructed tracks and down-sampled to $10\%$ of the total. For the electron sample, an overall relative improvement of $\sim 60\%$ is shown at $p_{\text{T}}<100$ MeV/$c$, with peaks of up to $\sim 80\%$ for the lowest momentum tracks in low pressure environments. This behavior is in agreement with Eqs. 32 and 33 which show a dependency of the $1/p_{\text{T}}$ resolution on $1/\sqrt{N}$ and $1/\sqrt{L_{\textrm{Arm}}}$ respectively: using the mirroring technique, more space points of the tracks are used, resulting in larger $N$ and $L_{\textrm{Arm}}$. The dependence on the number of points is shown more clearly in Fig. 16, where the histograms are color coded for total number of points in the track, including those only accessible through the mirroring technique. Tracks containing more than 800 points are excluded for easier legibility of the results. It is shown that for the longest tracks, relative improvements of up to 80% can be achieved.
@@ 3.3 Tests and results: high-pressure sample
In Fig. 17 we show the relative momentum resolution and bias as a function of the true momentum for the three particle types present in the sample. At lower momenta ($p_{\text{true}}<1\text{GeV/}c$) the resolution is close to 2% for pions and muons while it is closer to 8% for the protons. In this momentum region the multiple scattering component of the resolution is dominant. This component is inversely proportional to the particles’ $\beta$ factor. With a given momentum, the proton, having a higher mass, will always have a smaller $\beta$ and thus a worse expected resolution. Furthermore, while muons and protons at lower energies have the chance to produce longer and even looping tracks inside the detector, protons will tend to lose their energy more quickly, again due to their masses. This is clearly shown in Fig. 18(b) where the average particle lengths as a function of their true momentum is shown. Somewhat significant biases are also shown at these lower momenta, especially for protons. For momenta $p_{\text{true}}>1\leavevmode\nobreak\ \text{GeV/}c$, the resolution is comparable for the three particle types and increases slowly with the particle momentum. At higher momenta the point resolution component is dominant, so a direct proportionality on the momentum is expected, with no distinction between the particle types. An inverse dependency on the lever arm and number of points and thus indirectly on the length is also expected, but as shown in Fig. 18(b), in this momentum range the average length of the track becomes roughly the same for all particle types.
@@ Acknowledgments
Figure 22: Relative momentum resolution (a) and bias (b) as function of the tracks’ lever arm $L_{\textrm{Arm}}$ for the HP sample. The two properties are defined as $\mu$ and $\sigma$ of Gaussian fits of the momentum fractional residuals $p_{\text{reco}}/p_{\text{true}}-1$. The three particle types (protons, muons and pions) are drawn separately.